OpenStudy (anonymous):

The length of a rectangle is four more than twice the width. The perimeter of the rectangle is 44 meters. Find the length?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

k ! W+L = p Right ? P=44/2

OpenStudy (anonymous):

@xoya : ????

OpenStudy (anonymous):

And we have a system now : W + L = 22 L = 2W+4 K ?:)

OpenStudy (anonymous):

And we have system :) If have a problem send a massage :)

OpenStudy (anonymous):

@xoya : ???

OpenStudy (anonymous):

Not getting you sorry :/

OpenStudy (anonymous):

See we can just frame this question in an equation Like 2x+4

OpenStudy (anonymous):